# Evalet: Evaluating Large Language Models by Fragmenting Outputs into Functions

### v2

@@ 6.1.2. Conditions
Participants analyzed LLM outputs and their evaluations across two tasks in two conditions: Fragmented and Holistic (Fig. 6). The Fragmented condition was the full Evalet interface, without the holistic justifications (i.e., summaries of the fragment-level justifications for each output). The Holistic condition was a version of Evalet with only the holistic justifications. We chose this design as it closely resembles existing LLM-as-a-Judge approaches (Zheng et al., 2023; Kim et al., 2024a), but ensures that the evaluations in both conditions contain the same information. To ensure a fairer comparison, the Holistic condition also summarizes the holistic justification into a single label for each output (Fig. 6A), serving like a function description but for the whole output. These labels were embedded, clustered, and visualized the same way as in the Fragmented condition (Fig. 6C). For each output, this condition also highlights fragments relevant to each criterion (Fig. 6B), similar to prior work (Kim et al., 2024a), and allows users to flexibly select any fragments in the outputs to add as positive, negative, or excluded examples for a criterion.
This figure compares interface components between the Fragmented (treatment) and Holistic (control) study conditions. In the Details Tab, the Fragmented condition displays individual function label’s evaluations, while the Holistic condition shows only a summary of overall justification per output. In the Evaluation Details view, the Fragmented condition presents separate functions and reasoning for each function label, whereas the Holistic condition provides a overall justification and score. In the Map Visualization, the Holistic condition’s each point represents an entire output based on the overall evaluation embedding, instead of individual functions.
@@ 6.1.3. Tasks
Participants evaluated LLM outputs for the same two generation tasks: (1) writing a short horror story from a given set of keywords, and (2) writing an advertisement post for social media for a given product description. We chose these two tasks as they involve subjectivity, require no prior expertise, have similar input-output lengths, and have been explored by prior work (Kim et al., 2023a; Suh et al., 2024; Yuan et al., 2022). For each task, we created a dataset of 100 inputs and then generated outputs using gpt-4o-mini-2024-07-18, emulating a scenario of evaluating a relatively low-performing model. Then, we pre-evaluated these datasets using our approach to ensure that all participants, irrespective of condition, received the same evaluations for each task. Specifically, we used the following criteria for each task: (1) Horror Atmosphere for short horror stories (i.e., creating immersive and constant fear or psychological anxiety), and (2) Emotional Effect for the advertisement posts (i.e., effectively eliciting meaningful and genuine emotional responses from viewers). Since our participants were more fluent in Korean, we built the datasets in Korean, and added one line to our evaluation prompt to instruct the LLM to return function labels and justifications in Korean to minimize fluency-related effects. Full details on the datasets and criteria in Appendix C.
@@ 6.2. Results
This figure compares the Fragmented and Holistic conditions in two sections: output-related issues and evaluation-related issues. Each section includes a bar chart showing the average number of identified issues with standard deviation, and stacked bar charts showing the distribution of participant’s 7-point Likert ratings for importance, comprehensiveness, and actionability of the identified issues. Participants identified significantly more output-related issues in the Fragmented condition (M=3.40) than in the Holistic condition (M=2.30). Ratings also suggest that output-related issues in the Fragmented condition were seen as more actionable.
@@ 6.2.1. Verifying Evaluations
Participants identified significantly more issues with the LLM-based evaluations in the Fragmented condition (, , , ) (Fig.7). Participants mentioned that this was attributed to how it was easier to read and understand the fragment evaluations. While participants had to read entire outputs and overall justifications in the Holistic condition, they only had to review individual fragments, their function labels, and the accompanying shorter justifications in the Fragmented condition. P1 noted that “the range of text that I had to read was smaller so it was less time-consuming to interpret each data point.” As a result, P5 mentioned, “I read each evaluation more carefully and I was able to concentrate on each one more.”
@@
Furthermore, participants also mentioned how, as it was easier to verify each individual evaluation, it was also easier to verify their consistency. P3 explained, “As [each output’s] evaluation is split [into multiple fragment-level functions], I could see multiple evaluations [for fragments from different outputs] together and easily compare them, so I tended to focus on that.” In contrast, with the Holistic condition, P4 mentioned how “while I could see general trends in the evaluations, I could not directly compare them” due to the amount of text (i.e., outputs and overall justifications) to compare and reason about. Most participants (7/10) recognized the importance of identifying when the LLM-based evaluator is consistent or inconsistent, and explicitly focused on verifying this. P8 mentioned: “the problem [of inconsistent evaluations] is more important so no other issues really came into sight.”
@@
As participants’ ability to verify the evaluations differed in each condition, their trust and reliance on these evaluations also differed. As participants could more “confidently” (P1) verify evaluations in Fragmented, participants mentioned how they could judge their trust in the evaluations more carefully. P7 mentioned: “it’s not that I have more trust but instead that it is easier to verify my trust.” Three participants explained how they “empathized” with the LLM evaluator—explaining that they may not agree with its evaluation but understood why it returned such an evaluation. As a result, participants mentioned how they were able to develop more informed trust about the model by identifying the fragment-level functions for which they agreed with the LLM evaluator and when the evaluator is consistent or not. P5 explained: “I was wondering whether the evaluator had a bias when evaluating [a certain function] so I looked at these [clusters] more. My conclusion was that it seems like the LLM evaluator considers the aspect as negative most of the times, but there is a slight fluctuation.”
@@
On the other hand, as participants found it more cognitively demanding to verify evaluations in the Holistic condition, they also struggled to develop more informed trust. Some participants mentioned how they trusted the holistic evaluations despite not carefully inspecting them. For example, P7 mentioned that “if an output got a score of 100% and the [one-line justification summary] seems to make sense, I just move on”. Others mentioned not trusting the holistic evaluations at all as they could not confidently verify them: “I didn’t look at the summary or justification at all because I just didn’t have trust in them” (P5). Interestingly, regardless of their trust in the evaluations, most participants (7/10) mentioned relying on the evaluation scores to decide which outputs to explore—often focusing on extreme cases (i.e., scores of 100% or 0%)—even without fully understanding why those outputs received those scores.
@@ 6.2.2. Identifying Model Issues
Overall, participants identified a similar number of issues with the task LLM’s outputs in both conditions (, , , ). However, participants rated that they were significantly more confident that they could act on and resolve the issues identified with the Fragmented condition (, , , ) (Fig. 7).
According to participants, this result can be attributed to their trust and dependency in the LLM-based evaluations. As participants developed more informed trust regarding the evaluations in the Fragmented condition, they used the evaluations as guidance when inspecting the quality of the outputs. Participants followed the fragment-level evaluations to inspect outputs “piece-by-piece” (P1) and to inspect “each output more specifically” (P2). Furthermore, participants mentioned how the Fragmented condition allowed them to explore output issues from a “wider perspective” (P5) by exploring similar issues across outputs, or consider more “diverse characteristics” (P6) with regards to the criterion.
In contrast, in the Holistic condition, as participants could not adequately gauge their trust in the evaluations, they frequently mentioned not relying on the LLM-based evaluations and instead manually inspecting the outputs themselves. For example, P3 mentioned: “In the [Holistic condition], I had to read all of the [output] and also the justification, so I focused only on the [output] and tended to not look at the justification.” By manually reviewing the outputs themselves, participants not only lost efficiency benefits from the LLM-based evaluations but they also tended to focus on more abstract or surface-level issues regarding the model outputs (e.g., overall writing quality, coherency, logic). Besides being less concrete, P8 also explained how these broader issues could be more challenging to resolve: “The issues I identified seem more related to the limitations of the model itself […] No matter what feedback I give, it will be difficult to resolve these.” This is also reflected in the interaction logs, where participants in Holistic frequently viewed each output in detail (, , , ), while participants in Fragmented interacted more frequently with the Explore Tab and Map Visualization, selecting and navigating between data points (, , , ).
Although the Fragmented condition helped participants find more actionable issues, they also mentioned how it had limitations. Specifically, participants mentioned how they tended to lose sight of the “bigger picture” (P2)—including the overall structure, coherency, and context of the outputs. As a result, participants mentioned how they appreciated the Holistic condition as it allowed them to compare these holistic aspects of outputs. In fact, after exploring fragment-level functions, participants in the Fragmented condition frequently went back to the Database Tab as this was the only tab that allowed them to look at outputs one-by-one and compare them.
@@ 6.2.3. Correcting Evaluations
Participants found the Fragmented condition helpful for skimming through evaluations to identify potential examples for the criteria. However, in Fragmented, participants had to add the entire fragment that the system extracted as examples, which sometimes spanned multiple sentences. In contrast, the Holistic condition allowed them to manually select shorter fragments to add as examples, which they used to precisely select only the most relevant content. For example, P5 hesitated to add examples in the Fragmented condition: “since the whole [fragment] will be considered in future evaluations, I worry that [the evaluator] will interpret it differently.” Qualitative analysis of results indicated that the lower success rate in Fragmented for the horror story task stemmed from this limitation: automatically extracted fragments contained multiple sentences but participants likely added these as examples due to a short phrase within them. This suggests the need for a combined approach: fragment evaluations to identify evaluation issues, with the ability to select specific sub-fragments to precisely express intended corrections.
@@ 6.2.4. Exploration with Multiple Criteria
By exploring fragment-level functions for a new criterion (RQ4), participants were able to gain new insights about the outputs and the evaluations. For example, while comparing the evaluations for the same output on different criteria, P7 noticed how “the same fragment can receive completely opposite evaluations for different criteria” and that she wanted to “compare the distribution of evaluations in one cluster with those in a different criterion’s cluster.” P1 also mentioned how, as the new criterion focused on the overall content of the outputs rather than specific stylistic choices, she was able to “understand the outputs better”. She noted how there can be a “hierarchy” between criteria, where more general criteria should be evaluated first to understand the content of outputs, followed by more detailed and specific criteria. P6 explained how they could use these contrasting evaluations to decide on what outputs to use for “different use cases and applications”.
Participants noted that, by surfacing fragment-level functions relevant to each criterion instead of simply providing a score, Evalet allowed them to “more deeply understand and define each criterion” (P4). P2 mentioned: “When one doesn’t really know what is relevant to a criterion, they could just add an abstract description of the criterion [into the system], and see the LLM evaluations and clusters to learn more and concretize the criterion further.” P5 also reflected this sentiment: “[In practice], one needs to revise their evaluation criterion by actually evaluating outputs to see [how they match with the criterion], but it seems like this is already doing all of that for us.” Participants’ comments alluded to a process like inductive coding, where one starts with a broad and abstract criteria and, through the process of reviewing outputs, identifies relevant fine-grained functions that can concretize this criterion.
This figure presents the distribution of participants’ ratings for perceived workload across five NASA-TLX dimensions—Mental Demand, Temporal Demand, Effort, Frustration, and Performance—under Fragmented and Holistic conditions. Each bar shows a 7-point Likert scale distribution from Very Low (1, dark blue) to Very High (7, dark red). Overall, participants reported similar workload levels across both conditions, with generally high perceived workload due to the task demands.
@@ 6.2.5. Perceived Workload
Figure 8 shows that overall perceived workload was similar in both conditions (, , , ). This can be attributed to how each condition led to different distributions of cognitive effort. In Fragmented, verifying the evaluations required less effort which freed participants to compare and explore these evaluations. In contrast, Holistic led participants to expend most of their effort into manually reviewing outputs one by one. Also, while the fragment-level functions supported more fine-grained exploration and analysis, some participants found the number of functions to be overwhelming. For example, P7 mentioned that “The [Fragmented] visualization felt a bit complex and had too many colors, which made it hard to see what information I should focus on. In contrast, the [Holistic] visualization was much easier and didn’t feel tiring to look at.”
@@ 7. Example Cases
To demonstrate the generalizability of functional fragmentation and the insights that can be gained through it, we present three example cases of the approach with diverse LLMs and tasks. Specifically, we evaluate: (1) metacognition in reasoning LLMs, (2) harmlessness in user-LLM conversations, and (3) social intelligence in agent simulations. In this section, we briefly introduce the data that was evaluated and qualitative observations from the evaluations.
@@ 7.1. Metacognition in Reasoning
Reasoning-based LLMs, which generate explicit “reasoning” traces before providing final answers, have become increasingly prominent due to their advanced performance (Jaech et al., 2024). However, assessing these lengthy reasoning traces can be challenging. We applied functional fragmentation to the OpenThoughts dataset (Team, 2025), which contains the reasoning traces generated by DeepSeek-R1 (Guo et al., 2025) on various problems. Specifically, we sampled 210 data points (35 per dataset domain) and evaluated them on Metacognitive Insight (i.e., can the model actively reflect upon, regulate, and articulate their thought processes during reasoning).
As seen in Figure 9A, functional fragmentation surfaces diverse reasoning steps by the LLM that resemble human metacognition. For example, the fragment-level functions reveal self-questioning behaviors, explicit acknowledgment of uncertainties, and proactive consideration for edge cases. However, while evaluations tend to be overly positive, a closer look of functions shows that the evaluator credits metacognitive-like statements, even when they may not be beneficial. For example, in a reasoning trace, the model asks itself a question but will then immediately answer it, which superficially mimics human-like thought patterns but may not indicate authentic introspection or uncertainty. From here, practitioners could refine the evaluation criterion further to assess the actual impact of these metacognitive-like behaviors.
This figure presents a visual overview of clusters identified across three evaluation tasks: metacognitive insight in reasoning tasks, harmlessness in user-LLM conversations, and social intelligence in simulated interactions. Each subfigure displays a 2D embedding space of functions, where similar behaviors are grouped into color-coded clusters and annotated with representative functions that are discussed in the paper.

### v3

@@ 6.1.2. Conditions
Participants analyzed LLM outputs and their evaluations across two tasks in two conditions: Fragmented and Holistic (Fig. 6). The Fragmented condition was the full Evalet interface, without the holistic justifications (i.e., summaries of the fragment-level justifications for each output). The Holistic condition was a version of Evalet with only the holistic justifications, which closely resembles existing LLM-as-a-Judge approaches (Zheng et al., 2023; Kim et al., 2024b) but ensures that evaluations in both conditions contain the same information. To ensure a fairer comparison, the Holistic condition also summarizes the holistic justification into a single label for each output (Fig. 6A), serving like a function description but for the whole output. These labels were embedded, clustered, and visualized the same way as in the Fragmented condition (Fig. 6C). For each output, the Holistic condition also highlights fragments relevant to each criterion (Fig. 6B), similar to prior work (Kim et al., 2024b), and allows users to flexibly select any fragments in the outputs to add as positive, negative, or excluded examples for a criterion.
@@ 6.1.3. Tasks
Participants evaluated LLM outputs for the same two generation tasks: (1) writing a short horror story from a given set of keywords, and (2) writing an advertisement post for social media for a given product description. We chose these two tasks as they involve subjectivity, require no prior expertise, have similar input-output lengths, and have been explored by prior work (Kim et al., 2023b; Suh et al., 2024; Yuan et al., 2022). For each task, we created a dataset of 100 inputs and then generated outputs using gpt-4o-mini-2024-07-18, emulating a scenario of evaluating a relatively low-performing model. Then, we pre-evaluated these datasets using our approach to ensure that all participants, irrespective of condition, received the same evaluations for each task. Specifically, we used the following criteria for each task: (1) for short horror stories (i.e., creating immersive and constant fear or psychological anxiety), and (2) for the advertisement posts (i.e., effectively eliciting meaningful and genuine emotional responses from viewers). Since participants were more fluent in Korean, we built the datasets in Korean, and added one line to our evaluation prompt to instruct the LLM to return function labels and justifications in Korean to minimize fluency-related effects. Full details on the datasets and criteria in Appendix C.
@@ 6.2.1. Verifying Evaluations
Participants identified significantly more issues with the LLM-based evaluations in the Fragmented condition (, , , ) (Fig. 7). Participants attributed this to how it was easier to read and understand the fragment evaluations. Instead of reading entire outputs and overall justifications as in the Holistic condition, participants only had to review individual fragments, their function labels, and their shorter justifications in the Fragmented condition. P1 noted that “the range of text that I had to read was smaller so it was less time-consuming to interpret each data point.” As a result, P5 mentioned, “I read each evaluation more carefully and I was able to concentrate on each one more.”
@@
Furthermore, participants mentioned how, as it was easier to verify each individual evaluation, it was also easier to verify their consistency. P3 explained, “As [each output’s] evaluation is split [into multiple fragment-level functions], I could see multiple evaluations [for fragments from different outputs] together and easily compare them, so I tended to focus on that.” In contrast, with the Holistic condition, P4 mentioned how “while I could see general trends in the evaluations, I could not directly compare them” due to the amount of text (i.e., outputs and overall justifications) to compare and reason about. Most participants (7/10) recognized the importance of identifying when the LLM-based evaluator is inconsistent, and explicitly focused on verifying this.
@@
As participants’ ability to verify the evaluations differed in each condition, their trust and reliance on these evaluations also differed. As participants could more “confidently” (P1) verify evaluations in Fragmented, they mentioned how they could judge their trust in the evaluations more carefully. P7 mentioned: “it’s not that I have more trust but instead that it is easier to verify my trust.” Three participants explained how they “empathized” with the LLM evaluator—explaining that they may not agree with its evaluation but understood why it returned such an evaluation. As a result, participants mentioned how they were able to develop more informed trust about the LLM evaluator by identifying the fragment-level functions for which they agreed with the evaluator and when it is consistent or not. P5 explained: “I was wondering whether the evaluator had a bias when evaluating [a certain function] so I looked at these [clusters] more. My conclusion was that it seems like the LLM evaluator considers the aspect as negative most of the times, but there is a slight fluctuation.”
@@
On the other hand, participants struggled to develop more informed trust in the Holistic condition. Some participants mentioned how they trusted the holistic evaluations despite not carefully inspecting them. For example, P7 mentioned that “if an output got a score of 100% and the [one-line justification summary] seems to make sense, I just move on”. Others mentioned not trusting the holistic evaluations at all as they could not confidently verify them: “I didn’t look at the summary or justification at all because I just didn’t have trust in them” (P5). Interestingly, regardless of their trust in the evaluations, most participants (7/10) mentioned relying on the evaluation scores to decide which outputs to explore—often focusing on extreme cases (i.e., scores of 100% or 0%)—even without fully understanding why those outputs received those scores.
@@ 6.2.2. Identifying Model Issues
Overall, participants identified a similar number of issues in the task LLM’s outputs in both conditions (, , , ). However, participants rated that they were significantly more confident that they could act on and resolve the issues identified with the Fragmented condition (, , , ) (Fig. 7).
Participants attributed this to their trust and dependency in the LLM-based evaluations. As participants developed more informed trust regarding the evaluations in the Fragmented condition, they used the evaluations as guidance when determining the quality of the outputs—inspecting outputs “piece-by-piece” (P1) and “more specifically” (P2). Furthermore, participants mentioned how the Fragmented condition allowed them to explore output issues from a “wider perspective” (P5) by exploring similar issues across outputs and considering more “diverse characteristics” (P6) regarding the criterion.
In contrast, in the Holistic condition, participants could not adequately gauge their trust in the evaluations, so they would frequently manually inspected the outputs themselves without relying on the LLM-based evaluations. For example, P3 mentioned: “In the [Holistic condition], I had to read all of the [output] and also the justification, so I focused only on the [output] and tended to not look at the justification.” By manually reviewing the outputs themselves, participants not only lost efficiency benefits from the LLM-based evaluations but they also tended to focus on more abstract or surface-level issues regarding the model outputs (e.g., overall writing quality, coherency, logic). As they were less concrete, P8 explained that these broader issues could be more challenging to resolve: “The issues I identified seem more related to the limitations of the model itself […] No matter what feedback I give, it will be difficult to resolve these.” This is also reflected in the interaction logs, where participants in Holistic frequently viewed each output in detail (, , , ), while participants in Fragmented interacted more frequently with the Explore Tab and Map Visualization, selecting and navigating between data points (, , , ).
Although the Fragmented condition helped participants find more actionable issues, they also mentioned how it had limitations. Specifically, participants mentioned how they tended to lose sight of the “bigger picture” (P2), referring to the overall qualities of outputs (e.g., structure, coherency, and context). As a result, participants mentioned how they appreciated the Holistic condition as it allowed them to compare these holistic aspects of outputs. In fact, after exploring fragment-level functions, participants in the Fragmented condition frequently went back to the Database Tab as this was the only tab that allowed them to look at outputs one-by-one and compare them.
@@ 6.2.3. Correcting Evaluations
Participants found Fragmented helpful for skimming through evaluations to identify potential examples for the criteria. However, for each example, Fragmented required participants to add the entire fragment that the system extracted, which sometimes spanned multiple sentences. Due to this, P5 hesitated to add examples in the Fragmented condition: “since the whole [fragment] will be considered in future evaluations, I worry that [the evaluator] will interpret it differently. In contrast, Holistic allowed participants to manually select shorter fragments to add as examples, which they used to precisely select only the most relevant content.” Qualitative analysis of results indicated that the lower success rate in Fragmented for the horror story task stemmed from this limitation: automatically extracted fragments contained multiple sentences but participants likely added these as examples due to a short phrase within them. This suggests the need for a combined approach: fragment evaluations to identify evaluation issues, with the ability to select specific sub-fragments to precisely express intended corrections.
In Fragmented, participants would verify how adding examples to the criteria corrected and influenced the evaluations by rerunning the evaluations and using the Map Visualization to check how the evaluations changed. Several participants confirmed that they could see their feedback being incorporated. For instance, P9 mentioned “[newly added examples] seemed to be reflected” after observing that functions close to an example all adopted the same rating. This illustrates how the Fragmented condition facilitated participants’ iterative refinement of criteria by helping them identify functions to add as examples, steer the evaluations based on those examples, and verify the effects of this steering.
@@ 6.2.4. Exploration with Multiple Criteria
By exploring fragment-level functions for a new criterion (RQ4), participants were able to gain new insights about the outputs and the evaluations. For example, P7 observed that the fragment “experience a true ‘smart life’ with it” in an advertisement was positively rated for , but negative for as it only provides an abstract suggestion—highlighting the challenge of satisfying multiple criteria simultaneously and the opportunities for model refinement. As a result, P7 mentioned how he wanted to“compare the distribution of evaluations in one cluster with those in a different criterion’s cluster.” P6 explained how they could use these seemingly conflicting evaluations to decide on what outputs to use for “different use cases and applications”. P1 also noted that there was a “hierarchy” between the criteria that helped her gain both a broad and detailed understanding of outputs. The criterion (e.g., whether keywords are properly integrated in the horror story) helped her understand the broad alignment of the output’s content and its overall flow. Then, with , she could narrow down to sentence-level stylistic details (e.g., whether sentences effectively evoke a horror atmosphere).
Participants noted that, by surfacing fragment-level functions relevant to each criterion instead of simply providing a score, Evalet allowed them to “more deeply understand and define each criterion” (P4). P2 mentioned: “When one doesn’t really know what is relevant to a criterion, they could just add an abstract description of the criterion [into the system], and see the LLM evaluations and clusters to learn more and concretize the criterion further.” P5 also reflected this sentiment: “[In practice], one needs to revise their evaluation criterion by actually evaluating outputs to see [how they match with the criterion], but it seems like this is already doing all of that for us.” Participants compared the system to a process like inductive coding, where one starts with a broad and abstract criteria and, through the process of reviewing outputs, identifies relevant fine-grained functions that can concretize this criterion.
@@ 6.2.5. Perceived Workload
Figure 8 shows that overall perceived workload was similar in both conditions (, , , ), which is attributed to how each condition led to different distributions of cognitive effort. In Fragmented, verifying the evaluations required less effort which freed participants to compare and explore these evaluations. In contrast, Holistic led participants to expend most of their effort into manually reviewing outputs one by one. Also, while the fragment-level functions supported more fine-grained exploration and analysis, some participants found the number of functions to be overwhelming. For example, P7 mentioned that “The [Fragmented] visualization felt a bit complex and had too many colors, which made it hard to see what information I should focus on. In contrast, the [Holistic] visualization was much easier and didn’t feel tiring to look at.”
@@ 7. Example Cases
To demonstrate the generalizability of functional fragmentation and the insights that can be gained through it, we present three example cases of the approach with diverse LLMs and tasks. Specifically, we evaluate: (1) metacognition in reasoning LLMs, (2) harmlessness in user-LLM conversations, and (3) social intelligence in agent simulations. In this section, we briefly introduce the data that was evaluated and qualitative observations from the evaluations. In Appendix E, we include further details and an additional example on computer use agents.
@@ 7.1. Metacognition in Reasoning
Reasoning-based LLMs generate explicit “reasoning” traces before providing final answers—leading to significant performance gains (Jaech et al., 2024). However, assessing long reasoning traces can be hard. We applied functional fragmentation to 210 reasoning traces (Team, 2025) from DeepSeek-R1 (Guo et al., 2025) on the criterion (i.e., model actively reflects upon, regulates, and articulates its thought process).
As seen in Figure 9A, functional fragmentation surfaces diverse reasoning steps from the LLM’s traces that resemble human metacognition. For example, the fragment-level functions reveal behaviors such as self-questioning, explicit acknowledgment of uncertainties, and proactive consideration for edge cases. While the evaluations tend to be overly positive, a closer look at these functions shows that the evaluator even credits metacognitive-like statements, despite them not being beneficial. For example, in a reasoning trace, the model asks itself a question but will then immediately answer it, which superficially mimics human-like thought patterns but may not indicate authentic introspection or uncertainty. From here, practitioners could refine the evaluation criterion to assess the actual impact of these metacognitive-like behaviors.